# A Transfer Attack to Image Watermarks

### v2

@@ 5.1 Notations
We adopt $T$ to denote the target decoder and $\left\{S_{i}\right\}_{i=1}^{m}$ to represent the $m$ surrogate decoders. $w$ denotes the ground-truth watermark and $x_{w}$ denotes a watermarked image embedded with $w$. $\delta$ denotes a perturbation that satisfies the constraints of the optimization problem in Equation 3. For simplicity, we assume the target decoder and surrogate decoders use the same watermark length, which we denote as $n$. $T(\cdot)_{j}$ and ${S_{i}}(\cdot)_{j}$ denote the $j$th bit of the watermarks decoded by $T$ and $S_{i}$ for an image, respectively. Moreover, we consider inverse watermarks as the target watermarks in our transfer attack.
5.2 Formal Definitions and Assumptions
@@
To analyze the transferability between surrogate decoders and target decoder, we need to formally quantify their relationships. Towards this goal, we define unperturbed similarity, positive transferring similarity, and negative transferring similarity. Unperturbed similarity measures the probability that the watermarks decoded by the target decoder and surrogate decoders are identical for a watermarked image $x_{w}$ without perturbation, which is formally defined as follows:
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bit of the watermark decoded by $S_{i}$ for the perturbed image is the inverse of the $j$th bit of the watermark decoded by $S_{i}$ for the watermarked image with probability $q_{ij}$. Formally, we have the following:
@@
For any watermarked image $x_{w}$, the bits of the watermark decoded by $T$ are mutually independent and the probability that the $j$th bit of the decoded watermark matches with that of the ground-truth watermark $w$ is $\beta_{j}$, where $1\leq j\leq n$.
$\beta_{1},\beta_{2},\cdots,\beta_{n}$ can be estimated in experiments. Next, we introduce three assumptions regarding the surrogate decoders and the target decoder. For bit-level dependency, it is assumed that each bit of the watermark decoded by the target decoder only depends on the corresponding bits of the watermarks decoded by the surrogate decoders. Formally, we have the assumption as follows:
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, each bit of the watermark decoded by $T$ for the perturbed image only depends on the corresponding bits of the watermarks decoded by the surrogate decoders for the perturbed image. Formally, we have the following:
@@
$=Pr(T(x_{w}+\delta)_{j}\mid S_{1}(x_{w}+\delta)_{j},\cdots,S_{m}(x_{w}+\delta)_{j}),\forall j.$
Additionally, since all the surrogate decoders are trained independently with different subsets of data, we further assume that the watermarks decoded by the surrogate decoders are independent and conditionally independent given the watermark decoded by the target decoder for the watermarked or perturbed image.
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bits of the watermarks decoded by the $m$ surrogate decoders for the perturbed image are independent. Formally, we have the following:
@@
$=Pr(S_{1}(x_{w}+\delta)_{j})\cdots Pr(S_{m}(x_{w}+\delta)_{j}),\forall j.$
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bits of the watermarks decoded by the $m$ surrogate decoders for $x_{w}+\delta$ are independent when the $j$th bit of the watermark decoded by the target decoder for $x_{w}$ or $x_{w}+\delta$ is given. Formally, we have:
@@ 5.3 Deriving $Pr(T(x_{w}+\delta)_{j}=1-T(x_{w})_{j})$
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bit of the watermark decoded by $T$ for $x_{w}+\delta$ is the inverse of the $j$th bit of the watermark decoded by $T$ for $x_{w}$ with probability $a_{ij}k_{ij}+(1-a^{\prime}_{ij})(1-k^{\prime}_{ij})$, given that the $j$th bits of the watermarks for the $x_{w}+\delta$ and $x_{w}$ decoded by $S_{i}$ are inverse. Conversely, it occurs with probability $(1-b^{\prime}_{ij})k^{\prime}_{ij}+b_{ij}(1-k^{\prime}_{ij})$ given that the $j$th bits of the watermarks for $x_{w}+\delta$ and $x_{w}$ decoded by $S_{i}$ are identical. Formally, we have:
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bit of the watermark decoded by $T$ for the perturbed image is the inverse of the $j$th bit of the watermark decoded by $T$ for the watermarked image with probability $c_{ij}q_{ij}+c^{\prime}_{ij}(1-q_{ij})$. Formally, we have the following:
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bit of the watermark decoded by $S_{i}$ for the perturbed image is the inverse of the $j$th bit of the watermark decoded by $S_{i}$ for the watermarked image with probability $\frac{c_{ij}q_{ij}}{c_{ij}q_{ij}+c^{\prime}_{ij}(1-q_{ij})}$, given that the watermarks for the perturbed image and the watermarked image decoded by $T$ are inverse. Formally, we have the following:
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bit of the watermark decoded by $T$ for the perturbed image is the inverse of the $j$th bit of the watermark decoded by $T$ for the watermarked image with probability $p_{j}$, when the watermarks for the perturbed image decoded by the $m$ surrogate decoders are given. Formally, we have the following:
@@ 5.4 Deriving $Pr(T(x_{w}+\delta)_{j}=w_{j})$
The bitwise accuracy between the decoded watermark $T(x_{w}+\delta)$ and the ground-truth watermark $w$ is used to detect whether the perturbed image $x_{w}+\delta$ is AI-generated. Therefore, given the watermarks decoded by the $m$ surrogate decoders for the perturbed image $x_{w}+\delta$, we derive an upper bound and a lower bound of the probability that the watermark decoded by the target decoder matches with the ground-truth watermark after adding the perturbation. Formally, based on the theorems in Section 5.3, we further derive the following theorem:
Theorem 3.
Given any watermarked image $x_{w}$, $\delta$ satisfying the constraints of the optimization problem in Equation 3, and the watermarks decoded by the $m$ surrogate decoders for the perturbed image, we have an upper bound and a lower bound of the probability that the $j$th bit of the watermark decoded by $T$ for the perturbed image matches with the $j$th bit of the ground-truth watermark $w$ as follows:
$Pr(T(x_{w}+\delta)_{j}=w_{j}\mid S_{1}(x_{w}+\delta)_{j},\cdots,S_{m}(x_{w}+\delta)_{j})$ (15)
$\geq\mathop{max}(\beta_{j}-p_{j},0),$
$\leq 1-|p_{j}+\beta_{j}-1|.$
Such upper bound and lower bound quantify the transferability of our transfer attack. Note that we derive both an upper bound and a lower bound of the probability because we consider double-tail detectors.
6 Experiments
6.1 Experimental setup
Datasets:  In our experiments, we utilize AI-generated images for watermarking model training and testing. For the target watermarking models, we use two public datasets [21, 22] generated by Stable Diffusion and Midjourney respectively. Following HiDDeN [2], we randomly select 10,000 images from each dataset to train the target watermarking encoders and decoders. For testing, we randomly sample 1,000 images from the testing set of each dataset, embed the ground-truth watermark into each of them using a target encoder, and then find the perturbation to each watermarked image using different methods. To train the surrogate watermarking models, we sample 10,000 images from another public dataset [23] generated by DALL-E 2, i.e., the surrogate dataset consists of these 10,000 images. The input image size of the watermarking models is $128\times 128$ pixels.
Target and surrogate watermarking model architectures and watermark leng-
-ths:
  For surrogate watermarking models, we employ the convolution neural network architecture described in HiDDeN with the watermark length of 30 bits. We use the same architecture and watermark length for all surrogate watermarking models for two main purposes. Firstly, it can give us a clearer understanding of our attack’s transferability, especially when the target watermarking model and surrogate watermarking models have different architectures and watermark lengths. Secondly, it allows for a better analysis of the impact of the number of surrogate watermarking models on our attack’s transferability, without the confounding effects of variations in architecture and watermark length of the surrogate watermarking models.
For the target watermarking model, we use different choices of architectures and watermark lengths to analyze the transferability of our transfer attack in different scenarios. Specifically, we consider the following architectures for the target watermarking model.
CNN. The target watermarking model has convolution neural network architecture. In particular, the encoder consists of 4 convolutional blocks, while the decoder consists of 7 convolutional blocks. Each block integrates a Convolution layer, Batch Normalization, and ReLU activation.
ResNet. To further evaluate our transfer attack when the target watermarking model’s architecture is more sophisticated than that of the surrogate watermarking models, we also adopt ResNet for the target watermarking model. In particular, the encoder consists of 7 convolutional blocks and the decoder is the ResNet-18.
Common post-processing methods:  These methods can be applied in the no-box setting. Therefore, we compare our transfer attack with them. Specifically, we consider the following common post-processing methods.
Gaussian blur. This method smooths the image by averaging pixel values with their neighbors. It applies a Gaussian filter with kernel size of $k\times k$ to an image, characterized by a bell-shaped curve with a zero mean and a standard deviation of $\sigma$. A larger $\sigma$ causes more pronounced blurring, which results in lower watermark detection rate and image quality. Following Jiang et al. [11], we set $k=5$ and vary $\sigma$.
Brightness/Contrast. This method modifies the brightness and contrast of an image by adjusting pixel values throughout the image. Specifically, it operates by either increasing or decreasing these values to make the image brighter or darker. The process is regulated by two factors: a brightness factor $B$ and a contrast factor $C$. Formally, for each pixel value $v$, the method transforms it to $Cv+B$. Following Jiang et al. [11], we set $B=0.2$ and vary $C$.
(a) 20 bits
(b) 30 bits
(c) 64 bits
(a) 20 bits
(b) 30 bits
(c) 64 bits
(a) 20 bits
(b) 30 bits
(c) 64 bits
Transfer attacks:  We compare with the following attacks.
WEvade-B-S [11]. This method trains one surrogate watermarking model. During attack, PGD is employed to perturb a watermarked image such that the watermark decoded by the surrogate decoder for the perturbed image matches with a preset randomly generated watermark. In our experiments, we train the surrogate watermarking model on 10,000 images generated by DALL-E 2. To give advantages to this transfer attack, we assume the attacker knows the architecture of the target watermarking model and uses it as the architecture of the surrogate watermarking model.
AdvEmb-RN18 [12]. This method uses a ResNet-18 pretrained on ImageNet to generate a feature embedding for a watermarked image. Then it uses PGD to perturb the image such that the embedding of the perturbed image lies far from the one of the watermarked image.
AdvCls-Real&WM [12]. This method involves training a surrogate classifier using both watermarked and non-watermarked images. The watermarked images are generated by the target GenAI and are watermarked by the target watermarking model, and the non-watermarked images are drawn from a distribution different from the one of the images generated by the target GenAI. During attack, PGD is employed to perturb a watermarked image such that it is misclassified by the surrogate classifier as a non-watermarked image. In our experiments, we utilize images generated by DALL-E 2 as the non-watermarked images. More specifically, the training set for the surrogate classifier comprises 8,000 images generated by Stable Diffusion (or Midjourney) and watermarked by the target watermarking model, and 8,000 non-watermarked images generated by DALL-E 2. The surrogate classifier is based on the ResNet-18 architecture.
AdvCls-WM1&WM2 [12]. This method is the same as AdvCls-Real&WM except for the training data used for the surrogate classifier. It assumes that the attacker has access to watermarked images generated by the target watermarking model with two different watermarks. The surrogate classifier is trained to distinguish the watermarked images embedded with one watermark and those embedded with the other watermark. In our experiments, the training set for the surrogate classifier comprises 8,000 images watermarked by the target watermarking model with one watermark, and 8,000 images watermarked by the target watermarking model with the other watermark, where both watermarks are randomly picked.
Evaluation metrics:  We mainly use two metrics: evasion rate and average perturbation. Evasion rate is defined as the proportion of the perturbed images that evade the target watermark-based detector. Average perturbation is the perturbation added to a watermarked image measured by $\ell_{\infty}$-norm averaged across 1,000 watermarked images in the testing set. Following Jiang et al. [11], given that the pixel value range of [0, 255] is normalized to [-1, 1], we adjust the perturbation by dividing it by 2. This adjustment ensures that the perturbation is expressed as a fraction of the full pixel value range [0, 255].
Parameter settings:  We use adversarial training to train watermarking models since it enhances robustness. Specifically, for common post-processing methods, we consider the following range of parameters during adversarial training for a target watermarking model: $Q\in[10,99]$ for JPEG, $\sigma\in[0,0.1]$ for Gaussian noise, $\sigma\in[0,2.0]$ for Gaussian blur, and $C\in[1,3]$ for Brightness/Contrast. For surrogate watermarking models, we adopt a smaller range of parameters for some common post-processing methods during adversarial training to achieve weaker robustness than the target watermarking model as follows: $Q\in[50,99]$ for JPEG, $\sigma\in[0,0.1]$ for Gaussian noise, $\sigma\in[0,1.0]$ for Gaussian blur, and $C\in[1,3]$ for Brightness/Contrast.
By default, we set $max\_iter=5,000$, $r=0.25$, $\epsilon=0.2$, and $\alpha=0.1$ for our transfer attack. Unless otherwise mentioned, we use Inverse-Decode to select a target watermark for a surrogate decoder, and Ensemble-Optimization to find the perturbation. $\alpha$ is increased when the number of surrogate watermarking models increases in order to satisfy the constraints of our optimization problem within $5,000$ iterations. The detailed settings for $\alpha$ for different number of surrogate models are shown in Table 1 in Appendix. Moreover, we use $\ell_{2}$-distance as the distance metric $l(\cdot,\cdot)$ for two watermarks.
6.2 Experimental Results
Our transfer attack is successful:  Figure 1 and Figure 2 respectively show the evasion rate and average perturbation of our transfer attack as the number of surrogate models increases when the target model uses different neural network architectures and watermark lengths on the two datasets. First, we observe that our transfer attack is successful in evading watermark-based detection. The evasion rate is higher than 85% and the average perturbation is no larger than 0.25 when 100 surrogate models are used, no matter what architectures and watermark lengths are used by the target model.
Second, when the target model’s architecture is more complex than the surrogate models’ architecture, our attack requires more surrogate models and larger average perturbation to succeed. For example, our attack uses 40 surrogate models and average perturbation 0.09 to achieve an evasion rate around 100% when the target model is CNN and uses 30-bit watermarks on Stable Diffusion, while our attack uses 60 surrogate models and average perturbation 0.14 to achieve a similar evasion rate when the target model is ResNet and uses 30-bit watermarks. Third, when the target model uses a longer/shorter watermark than the surrogate models, our attack requires more surrogate models and larger average perturbation to succeed. For example, our attack uses 40 surrogate models and average perturbation 0.08 to achieve an evasion rate around 85% when the target model is ResNet and uses 30-bit watermarks, while our attack uses 100 surrogate models and average perturbation 0.15 to achieve a similar evasion rate when the target model uses 64-bit watermarks.
Compare with common post-processing:  Figure 3 compares the average perturbations of common post-processing methods and our transfer attack when they achieve the same evasion rates for different target models on the two datasets. Given an evasion rate achieved by our transfer attack using a certain number of surrogate models, we set the parameter of a common post-processing method such that it achieves very close evasion rate. We observe that our transfer attack adds much smaller perturbations than common post-processing methods when achieving the same evasion rate. Note that for some target models, the Brightness/Contrast method can only achieve about 75% evasion rate at most. Thus, the curves for Brightness/Contrast are shorter in the corresponding graphs.
To get a more comprehensive comparison with common post-processing methods, Figure 10 and Figure 11 in Appendix show the comparison results using average perturbation measured by $\ell_{2}$-norm and average SSIM, respectively. SSIM measures the similarity between a perturbed image and the corresponding watermarked image by considering changes in texture, luminance, and contrast. Average SSIM is the SSIM averaged across 1,000 watermarked images and their perturbed versions in the testing set. We still observe that our transfer attack adds much smaller perturbations and thus better maintains the images’ visual quality than common post-processing methods when $\ell_{2}$-norm or SSIM is used to measure the perturbation.
Figure 4: Comparing evasion rates of existing and our transfer attacks when the target model is ResNet and uses watermarks with different lengths. Dataset is Stable Diffusion. Similar results for Midjourney are shown in Figure 12 in Appendix.
Compare with existing transfer attacks:  Figure 4 compares the evasion rates of existing and our transfer attacks when the target model is ResNet and uses watermarks with different lengths. We note that since AdvCls-Real&Wm and AdvCls-WM1&WM2 can only achieve average perturbation around 0.1 no matter how large the perturbation budget $r$ we set, we add one extra variant of our method with perturbation budget $r=0.1$ for comparison. For other transfer attacks, the perturbation budget $r$ is set to be 0.25. We observe that our transfer attack is much more effective than other existing transfer attacks. First, we observe that existing transfer attacks achieve almost 0 evasion rates except that AdvCls-WM1&WM2 achieves around 10% evasion rates on the Stable Diffusion dataset when the target model’s watermark length is 20 or 30 bits. This is the reason that existing works concluded that watermarking is robust in the no-box setting. However, our transfer attack achieves much higher evasion rates especially when the perturbation budget is large enough (e.g., $r=0.25$). We note that the average SSIM of our transfer attack is higher than 0.97 for $r=0.1$ and higher than 0.92 for $r=0.25$, which means that the image quality of the perturbed images is maintained well.
(a) Evasion rate
(b) $\ell_{\infty}$-norm perturbation
Figure 5: Evasion rate and average $\ell_{\infty}$-norm perturbation of our transfer attack when using different $\epsilon$.
Impact of $\epsilon$:  Figure 5 shows the evasion rate and average $\ell_{\infty}$-norm perturbation when our transfer attack uses different $\epsilon$ on Stable Diffusion dataset, where the target model is CNN and uses 30-bit watermarks. Figure 13 in Appendix shows the results on average $\ell_{2}$-norm perturbation and average SSIM. We observe that our attack requires less surrogate models to achieve the same evasion rate when $\epsilon$ is smaller. However, when $\epsilon$ is too small (e.g., 0.1), the evasion rate first increases and then decreases as the number of surrogate models increases. This is because the bitwise accuracy between the decoded watermark of the perturbed image and the ground-truth watermark becomes too small when $\epsilon$ is too small, and thus the perturbed image is still detected as AI-generated by the double-tail detector. Moreover, we observe that, for different $\epsilon$, the average perturbation (measured by $\ell_{\infty}$-norm, $\ell_{2}$-norm, or SSIM) keeps almost the same when achieving the same evasion rate. This is because the average perturbation and the evasion rate have strong correlations in our attack.
(a) Evasion rate
(b) $\ell_{\infty}$-norm perturbation
Figure 6: Evasion rates and average $\ell_{\infty}$-norm perturbations of different variants of our transfer attack.
Different variants of our transfer attack:  Our attack has two steps, each of which has multiple design choices. We use RD, RS, and ID to respectively denote Random-Different, Random-Same, and Inverse-Decode in the first step; and PA and EO to respectively denote Post-Aggregate and Ensemble-Optimization in the second step. We concatenate the symbols to represent a variant, e.g., RD-PA means using RD and PA as the two steps. Moreover, when PA is used, we further use suffix ”Mean” or ”Median” to indicate the adopted aggregation rule. For instance, RD-PA-Mean indicates that Mean is used as the aggregation rule.
Figure 6 shows the evasion rate and average $\ell_{\infty}$-norm perturbation when different variants of our transfer attack are used on Stable Diffusion dataset, where the target model is CNN and uses 30-bit watermarks. Figure 13 in Appendix shows the results on average $\ell_{2}$-norm perturbation and average SSIM. We observe that all the three variants using EO can evade detection successfully while none of the variants using PA can evade detection, which indicates that EO is much more effective at aggregating surrogate decoders than PA. Additionally, within the three variants using EO, RD and RS have similar results for all metrics and are less effective than ID. Specifically, ID-EO achieves 100% evasion rate with less number of surrogate models compared to RD-EO and RS-EO. For the same 100% evasion rate, ID-EO achieves a much lower average $\ell_{\infty}$-norm and $\ell_{2}$-norm perturbation and a much higher average SSIM than RD-EO and RS-EO.
(a) CNN
(b) ResNet
Figure 7: Comparing empirical bitwise accuracy and theoretical bounds of our transfer attack when the target model is (a) CNN and (b) ResNet.
Theoretical vs. empirical results:  Figure 7 shows the empirical average bitwise accuracy between the watermark decoded by the target decoder for a perturbed image and the ground-truth one, and our theoretical upper and lower bounds on Stable Diffusion dataset. The theoretical bounds are estimated as follows: we first measure the unperturbed similarity, transferring similarity, and attacking strength using the 1,000 testing images; then we compute the probability $p_{j}$ for each $j$th bit based on Theorem 2, following which we compute the upper and lower bounds of the probability for each bit of the watermark decoded by the target decoder to match with that of the ground-truth watermark; and finally we average these bounds across the bits. We observe that our theoretical upper and lower bounds match well with the empirical bitwise accuracy. Specifically, the empirical results fall within our theoretical bounds except a few cases. Such exception cases happen because the theoretical bounds are estimated using only 1,000 images, resulting in inaccurate estimations.
7 Discussion and Limitations
Figure 8: Evasion rate of our transfer attack when the target model is non-learning-based.
We focus on transfer attacks to learning-based target watermarking models because learning-based models are more robust than non-learning-based ones. For instance, common post-processing like JPEG compression can already remove watermarks embedded by a non-learning-based model, but learning-based models have good robustness against common post-processing [11]. However, our attack can also be applied to non-learning-based models. For instance, Figure 8 shows the evasion rate of our transfer attack when the target model is the watermarking model used by Stable Diffusion [24] as the number of surrogate models increases on the two datasets, where the surrogate models are the same as those in Section 6. Our results show that our transfer attack can still achieve high evasion rates when the target model is non-learning-based. Compared to when the target model is learning-based, our transfer attack requires more surrogate models to achieve high (e.g., 100%) evasion rates. This is because non-learning-based watermarking model is substantially different from learning-based, and thus our attack needs more surrogate models to increase diversity and thus enhance transferability.
8 Conclusion and Future Work
In this work, we find that watermark-based detection of AI-generated images is not robust to transfer attacks in the no-box setting. Given a watermarked image, an attacker can remove the watermark by adding a perturbation to it, where the perturbation can be found by ensembling multiple surrogate watermarking models. Our results show that transfer attack based on surrogate watermarking models outperforms those based on surrogate classifiers that treat a watermark-based detector as a conventional classifier. Moreover, leveraging surrogate watermarking models enables us to perform a rigorous analysis on the transferability of our attack. Interesting future work includes extending our transfer attack to text and audio watermarks, as well as designing more robust, especially certifiably robust, watermarks.
Appendix A Proof of Lemma 1

### v4

@@ B.1 Notations
We adopt $T$ to denote the target decoder and $\left\{S_{i}\right\}_{i=1}^{m}$ to represent the $m$ surrogate decoders. $w$ denotes the ground-truth watermark and $x_{w}$ denotes a watermarked image embedded with $w$. $\delta$ denotes a perturbation that satisfies the constraints of the optimization problem in Equation 3. We use $n_{s}$ and $n_{t}$ to denote the watermark length of surrogate decoders and target decoder respectively, while we define $n=\mathop{min}(n_{s},n_{t})$. $T(\cdot)_{j}$ and ${S_{i}}(\cdot)_{j}$ denote the $j$th bit of the watermarks decoded by $T$ and $S_{i}$ for an image, respectively. Moreover, we consider inverse watermarks as the target watermarks in our transfer attack.
B.2 Formal Definitions
@@
In this section, we will introduce the details of our formal definitions. We formally define unperturbed similarity, positive transferring similarity, negative transferring similarity, $q$-attacking strength, $\beta$-accurate watermarking, Bit-level dependency, Independency, and Conditional independency. Specifically, unperturbed similarity measures the probability that the watermarks decoded by the target decoder and surrogate decoders are identical for a watermarked image $x_{w}$ without perturbation, which is formally defined as follows:
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bit of the watermark decoded by $S_{i}$ for the perturbed image is the inverse of the $j$th bit of the watermark decoded by $S_{i}$ for the watermarked image with probability $q_{ij}$. Formally, for $i=1,2,\cdots,m$ and $j=1,2,\cdots,n_{s}$, we have the following:
@@
For any watermarked image $x_{w}$, the bits of the watermark decoded by $T$ are mutually independent and the probability that the $j$th bit of the decoded watermark matches with that of the ground-truth watermark $w$ is $\beta_{j}$, where $1\leq j\leq n_{t}$.
$\beta_{1},\beta_{2},\cdots,\beta_{n_{t}}$ can be estimated in experiments. Next, we introduce three assumptions regarding the surrogate decoders and the target decoder. For bit-level dependency, it is assumed that each bit of the watermark decoded by the target decoder only depends on the corresponding bits of the watermarks decoded by the surrogate decoders. Formally, we have the assumption as follows:
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, each bit of the watermark decoded by $T$ for the perturbed image only depends on the corresponding bits of the watermarks decoded by the surrogate decoders for the perturbed image. Formally, for $j=1,2,\cdots,n$, we have the following:
@@
$=Pr(T(x_{w}+\delta)_{j}\mid S_{1}(x_{w}+\delta)_{j},\cdots,S_{m}(x_{w}+\delta)_{j}).$
Additionally, since all the surrogate decoders are trained independently with different subsets of data, we further assume that the watermarks decoded by the surrogate decoders are independent and conditionally independent given the watermark decoded by the target decoder for the watermarked or perturbed image. We verify the independency assumption and the results are shown in Figure 14.
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bits of the watermarks decoded by the $m$ surrogate decoders for the perturbed image are independent. Formally, for $j=1,2,\cdots,n_{s}$, we have the following:
@@
$=Pr(S_{1}(x_{w}+\delta)_{j})\cdots Pr(S_{m}(x_{w}+\delta)_{j}).$
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bits of the watermarks decoded by the $m$ surrogate decoders for $x_{w}+\delta$ are independent when the $j$th bit of the watermark decoded by the target decoder for $x_{w}$ or $x_{w}+\delta$ is given. Formally, for $j=1,2,\cdots,n$, we have:
@@ B.3 Deriving $Pr(T(x_{w}+\delta)_{j}=1-T(x_{w})_{j})$
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bit of the watermark decoded by $T$ for $x_{w}+\delta$ is the inverse of the $j$th bit of the watermark decoded by $T$ for $x_{w}$ with probability $a_{ij}k_{ij}+(1-a^{\prime}_{ij})(1-k^{\prime}_{ij})$, given that the $j$th bits of the watermarks for the $x_{w}+\delta$ and $x_{w}$ decoded by $S_{i}$ are inverse. Conversely, it occurs with probability $(1-b^{\prime}_{ij})k^{\prime}_{ij}+b_{ij}(1-k^{\prime}_{ij})$ given that the $j$th bits of the watermarks for $x_{w}+\delta$ and $x_{w}$ decoded by $S_{i}$ are identical. Formally, for $j=1,2,\cdots,n$, we have:
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bit of the watermark decoded by $T$ for the perturbed image is the inverse of the $j$th bit of the watermark decoded by $T$ for the watermarked image with probability $c_{ij}q_{ij}+c^{\prime}_{ij}(1-q_{ij})$. Formally, for $j=1,2,\cdots,n$, we have the following:
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bit of the watermark decoded by $S_{i}$ for the perturbed image is the inverse of the $j$th bit of the watermark decoded by $S_{i}$ for the watermarked image with probability $\frac{c_{ij}q_{ij}}{c_{ij}q_{ij}+c^{\prime}_{ij}(1-q_{ij})}$, given that the watermarks for the perturbed image and the watermarked image decoded by $T$ are inverse. Formally, for $j=1,2,\cdots,n$, we have the following:
@@
For any watermarked image $x_{w}$ and $\delta$ satisfying the constraints of the optimization problem in Equation 3, the $j$th bit of the watermark decoded by $T$ for the perturbed image is the inverse of the $j$th bit of the watermark decoded by $T$ for the watermarked image with probability $p_{j}$, when the watermarks for the perturbed image decoded by the $m$ surrogate decoders are given. Formally, for $j=1,2,\cdots,n$, we have the following:
@@ B.4 Deriving $Pr(T(x_{w}+\delta)_{j}=w_{j})$
The bitwise accuracy between the decoded watermark $T(x_{w}+\delta)$ and the ground-truth watermark $w$ is used to detect whether the perturbed image $x_{w}+\delta$ is AI-generated. Therefore, given the watermarks decoded by the $m$ surrogate decoders for the perturbed image $x_{w}+\delta$, we derive an upper bound and a lower bound of the probability that the watermark decoded by the target decoder matches with the ground-truth watermark after adding the perturbation. Note that we consider $1\leq j\leq n_{t}$ in this section. Formally, based on the theorems in Section B.3, we derive the Theorem 1 in Section 5.
Our theoretical analysis demonstrates that the probability of the $j$th bit of the watermark decoded by $T$ for the perturbed image matching with the $j$th bit of the ground-truth watermark can be bounded, which can be used to compute the upper bound and lower bound of the expected bitwise accuracy between $T(x_{w}+\delta)$ and $w$. If such upper bound and lower bound fall between $\tau$ and $1-\tau$, the perturbed image is expected to evade the target watermark-based detector.
Appendix C Proof of Lemma 1